Solve
4x^2 - y^2 = 4
4x^2 + y^2 = 4

Answer by checkley71(8403)

4x~2-y~2=4
4x~2+y~2=4 adding these two equations we get
----------
8x~2 =8 or x~2=8/8 or x~2=1 or x=1 then
4*1~2+y=4 or 4+y=4 y=0